Let a,b,c be distinct non zero real numbers such that:
1-a³/a=1-b³/b=1-c³/c.Then value of a³+b³+c³ is?

Grade:8

1 Answers

Arun
5 years ago
Dear Divyashu
 
1 – a^3 = ka
 
a^3 = 1 + ka
 
similarily
 
b^3 = 1 + kb
 
c^3 = 1 + kc
 
hence 
a^3 + b^3 + c^3 = 3 + k (a + b +c)